### v1

# IFEvalCode: Controlled Code Generation

Abstract
Code large language models (Code LLMs) have achieved significant advancements in various code-related tasks, particularly in code generation, where the code LLMs produce the target code from natural language descriptions. However, in realistic scenarios, users often expect the return code to strictly follow the given detailed requirements in many aspects (e.g. the style of code, the number of code lines, or the number of lines), instead of only requiring the correctness of the generated code. Controlled code generation means that the generated response from code LLMs should adhere to specific human guidelines or standards, whereas the LLM should have a strong instruction-following capability in the field of the code. In this paper, we propose forward constraints generation and backward constraints generation for controlled code generation to enhance the capability of LLM in following human instructions. Then, we build a multilingual benchmark IFEvalCode to evaluate the code instruction-following capability of the LLMs. IFEvalCode consists of 1.6K samples (Python, Java, Javascript, Typescript, Shell, Cpp, and C-sharp) and each test sample contains the Chinese and English query. Different from the existing code benchmarks, we separately design the test function to verify the correctness of the code (Corr.) and whether the generated code follows the human instruction (Instr.). Extensive experimental results on IFEvalCode of 40+ LLMs emphasize that closed-source LLMs still dominate in controllable code generation compared to open-source LLMs and the ability of LLMs to generate controllable code is far behind its ability to generate correct code [^1] The code and dataset will be released.
1 Introduction
Large Language Models (LLMs) have made significant strides in code-related tasks, particularly in producing target code snippets from the given human description. It has captured considerable attention from academia and industry, owing to its practical relevance in software development. The current top-tier closed-source LLMs, such as GPT4.5 (gpt45, ) and Clude3.7 (claude37, ), can correctly generate the file-level code snippet most of the time. Besides, a series of open-source code LLMs, such as StarCoder1/2 (starcoder, ; starcoder2, ), Deepseek-Coder (deepseek_coder, ), OpenCoder (opencoder, ), and Qwen2.5-Coder (qwen_coder, ), has also become a force that cannot be ignored by industry and academia of code intelligence.
Most existing code benchmarks (mceval, ), such as HumanEval (codex, ), BigCodeBench (bigcodebench, ), and MBPP (mbpp, ) are designed to evaluate the correctness of generated code by code execution with the unit tests. Further, some benchmarks, such as Chatbot Arena (llm_as_a_judge, ) and CodeArena (codearena, ) are proposed to evaluate the alignment between the model-generated response and human preference using the LLM-as-a-Judge. For the general domain, IFEval (ifeval, ) is proposed to evaluate the proficiency of LLM in controlled text generation, where the instructions are amenable to objective verification of compliance. In the field of code, in addition to requiring the generation of correct code, users often also require the generation of code meeting various objective requirements, such as code style, variable naming, specific algorithms, time complexity, etc. Therefore, we try to explore the proficiency of LLMs in controlled code generation by designing a framework to enhance and evaluate the capabilities of in code instruction following.
Figure 1: Massively multilingual evaluation task comprised of three tasks, including code generation, code completion, and code explanation.
To evaluate the capability of generating the controlled code, we construct a multilingual benchmark using human annotators, IFEvalCode, comprising nearly 1.6K samples across 888 programming languages (Python, Java, CPP, C-Sharp, TypeScript, JavaScript, PHP, Shell), with each sample including both Chinese and English queries. Unlike existing benchmarks, IFEvalCode uniquely assesses two critical aspects: the correctness (code correctness) of the generated code and its adherence to human instructions (code controlability). In this paper, we introduce a novel framework combining forward constraints generation and backward generation for controlled code generation, aimed at enhancing the ability of LLM to follow human instructions in controlled code generation.
The contributions are summarized as follows:
  • We introduce IFEvalCode, a multilingual benchmark comprising 1,620 samples across 888 programming languages (Python, Java, JavaScript, TypeScript, Shell, C++, and C#) with each sample featuring paired Chinese and English queries to facilitate comprehensive evaluation in controlled code generation.
  • We integrate forward constraints generation and backward constraints generation to generate the instruction corpus IFEvalCode-Instruct, designed to enhance the capability of LLMs in adhering to human instructions during code generation. By jointly training to improve the controllability, accuracy, and alignment of generated code with specific user requirements, thereby advancing the performance of LLMs in structured and instruction-driven coding tasks.
  • We perform a comprehensive evaluation and analysis of existing 40+ LLMs on IFEvalCode, and several insightful findings are as follows: (1) Controlled code generation is challenging for LLMs. LLMs still have huge room for improvement and a long journey ahead to reach their full potential. (2) Closed-source models (e.g. Claude3.7) and LLMs with large sizes still significantly outperform open-source LLMs and LLMs with small sizes in controllable code generation. (3) The ability to generate specified code is significantly weaker than the ability to generate correct code.
2 Multilingual Controlled Code Generation
2.1 Task Definition
Given the $k$-th programming language $L_{k}\in\{L_{k}\}_{k=1}^{K}$ (in this work, $K=8$). We provide the problem description $q^{L_{k}}$ and instruction constraints $c^{L_{k}}=\{c^{L_{k}}_{1},\dots,c^{L_{k}}_{m}\}$ as input to the code LLMs $\mathcal{M}$ (Note: the constraints for code generation is language-specific). The LLMs generates the corresponding code $a^{L_{k}}$ by sampling from the code generation distribution $P(a^{L_{k}}|q^{L_{k}},c^{L_{k}};\mathcal{M})$. The generated code is then executed with the provided test cases $u_{L^{k}}$ to verify correctness. Specially, we design another check function $v^{L_{k}}$ for each query to determine whether the generated code is controllable meeting human instruction constraints. This process ensures that the generated code $a^{L_{k}}$ functionally aligns with the problem requirements and :
$r^{L_{k}}_{u}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}},e^{L_{k}};\mathcal{M});u^{L_{k}});\ \ \ \ \ \ r^{L_{k}}_{v}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}},e^{L_{k}};\mathcal{M});v^{L_{k}});$ (1)
where $\mathbb{I}(\cdot)$ is the code execution in a sandbox environment with the given test cases $u^{L_{k}}$ or the check function $v^{L_{k}}$. when the generated code $a^{L_{k}}$ passes the unit tests $u^{L_{k}}$, the evaluation result $r^{L_{k}}_{u}=1$, else $r^{L_{k}}_{u}=0$. For controlled code generation, the check function $v^{L_{k}}$ is used to check whether the generated code satisfy the human instruction. $u^{L_{k}}$ and $v^{L_{k}}$ separately denote the scores of code correctness and code controllability.
2.2 Overview of IFEvalCode
Table 1: Benchmark statistics.
Statistics Number
Problems $/810$ 810
Python $/206$ 206
Java $/204$ 204
Cpp $/200$ 200
Typescript $/200$ 200
Javascript $/200$ 200
Php $/200$ 200
Shell $/200$ 200
#Instruction Constraints
- min/max/avg length 2/15/5
Length
Chinese Question
min/max/avg 142142142/1,3731,3731 , 373/358358358 tokens
English Question
min/max/avg 135135135/1,3041,3041 , 304/340340340 tokens
In Table 1, the IFEvalCode consists of 1.6K problems. Each task contains about 1.3K questions, with more than 60 problems in each language. Each sample in IFEvalCode includes (en_question, zh_question, check_instruction, check_correctness), where ‘check_instruction‘ is used to check the code correctness and ‘check_correctness‘ judges whether the generated code follows the human instruction. We calculate the length of the question using the Qwen2.5-Coder tokenizer (qwen_coder, ) and count the number of instruction constraints for each sample. Each question contains 3 constraints and 100 tokens on average, where each prompt contains a problem description and the corresponding function or class signature.
In Table 2, we compare IFEvalCode with other code evaluation benchmarks in 6 aspects. Considering the evaluation cost and effectiveness, IFEvalCode has 1620 samples for 8 programming languages and 2 human languages. Specially, we introduce the two part of unit tests to check the code correctness and whether the code follows the human instruction.
Table 2: Comparison between IFEvalCode and other code benchmarks. Different from the existing benchmarks, IFEvalCode effectively provides a new comprehensive view of controlled code generation by adding the ‘check_instruction‘ part into the test samples.
Benchmark #Languages #Domains Size (Zh/En) Difficulty Human Check Check Code Controllability
HumanEval (codex, ) Python 3 164 (En) [Uncaptioned image]
MBPP (mbpp, ) Python Algorithm 974 (En) [Uncaptioned image]
BigCodeBench (bigcodebench, ) Python 7 2000+ (En) [Uncaptioned image] [Uncaptioned image]
CodeArena (codearena, ) 44 7 397+ (En) [Uncaptioned image] [Uncaptioned image]
FullStackBench (fullstackbench, ) 16 11 3374 (En) [Uncaptioned image] [Uncaptioned image]
LiveCodeBench (livecodebench, ) Python 3 450+ (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MultiPL-E (multipl_e, ) 19 3 12,667 (En) [Uncaptioned image] [Uncaptioned image]
McEval (mceval, ) 20 10+ 16K (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MdEval (mdeval, ) 40 10+ 39K (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DS-1000 (ds1000, ) Python Data Size 1000 (En) [Uncaptioned image] [Uncaptioned image]
IFEvalCode (Ours) 8 3 810/810 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(a) Domain types.
(b) Types of instruction constraints.
Figure 2: Figure 2(a) shows domain types and Figure 2(b) lists instruction types.
2.3 IFEvalCode Construction & Quality Control
To create IFEvalCode for controlled code generation, we propose a systematic human annotation and check procedure, which is guided by defined guidelines to guarantee accuracy and consistency. Figure 4 illustrates the overall process of dataset construction. We begin by collecting code snippets and code-related documents from websites and recruit 6 computer science graduates as annotators. The annotators must follow the provided guidelines (1) ensure the diversity of the questions. (2) the question and instruction constraints should be challenging for existing LLMs. (3) translate the English question into the Chinese question. To increase the difficulty of IFEvalCode, the annotator filter out the questions, which half LLMs (GPT4o, DeepSeek-V3, Claude3.7, etc) can correctly answer.
Figure 3: Examples of the verifiable instructions with ‘check_instruction’ in IFEvalCode.
Figure 4: Construction of the benchmark IFEvalCode and instruction corpus IFEvalCode-Instruct.
2.4 Reject-sampling Fine-tuning for Controlled Code Generation
Creating Queries and Unit Tests with Constraints
To create the instruction corpora, we collect the code-related websites and remove the HTML tags to obtain the code snippet with the corresponding natural language description as the seed data. Then, we select the select seed data from 8 languages and force the code LLM (Qwen2.5-Coder 32B) to generate a challenging question $q^{L_{k}}$ of programming language $L_{k}$. To verify the correctness of the response, we prompt the LLM to generate the test cases $u_{k}$ to create ‘check_correctness’. To independently check whether the model-generated response satisfies the constraints in the question, we assume that the response is correct and only check whether the generation code follows the user instructions. Therefore, for each query $q^{L_{k}}$ with constraints $c^{L_{k}}$, we create the unit tests $r_{u}^{L_{k}}$ to check the correctness of the response and $r_{u}^{L_{k}}$ to check the whether the generated code is under control.
Forward Constraints Generation
Given the recalled code-related documents from Common Crawl, we adopt Qwen2.5-Coder-32B to create new questions by drawing inspiration from the code-related documents for a general question. To effectively inject the additional instruction constraints into the question, we select the constraint types from the constraints types pool $C=\{c_{i=1}\}_{i=1}^{m}$ and force the model to generate a new question $q^{L_{k}}$ with constraints. After generating the question $q^{L_{k}}$ with constraints $c^{L_{k}}$, we force the LLM to generate $r^{L_{k}}_{u}$ and $r^{L_{k}}_{v}$ and the only keep the samples passing both unit tests. The generated corpus is denoted as $D_{s_{1}}$
Backward Constraints Generation
Since the forward constraints generation must first pre-define the constraint types and then generate the corresponding question, we introduce the backward constraints generation to synthesize the question with random instruction constraints. First, we prompt the LLM to generate the question $q^{L_{k}}$ and answer $a^{L_{k}}$. Then, we prompt the LLM to summarize which constraints the generated code complies with and obtain othe constraints $c^{L_{k}}$. Finally, we generate the unit tests $r^{L_{k}}_{u}$ and $r^{L_{k}}_{v}$, which are further used to verify the response and the created instruction corpus from backward constraints generation is denoted as $D_{s_{2}}^{L_{k}}$.
Joint Multilingual Rejecting Fine-tuning for Controlled Code Generation
Given the large-scale instruction corpus $D^{L_{k}}=D_{s_{1}}^{L_{k}}\bigcup D_{s_{2}}^{L_{k}}$ of different programming languages, the LLM is jointly optimized for response generation and unit tests generation of different programming languages:
$\mathcal{L}_{all}=-\sum_{k=1}^{K}\mathbb{E}_{D^{L_{k}}}\log P(a^{L_{k}}|q^{L_{k}},c^{L_{k}})-\sum_{k=1}^{K}\mathbb{E}_{D^{L_{k}}}\log P(u^{L_{k}}_{v},u^{L_{k}}_{r}|q^{L_{k}},c^{L_{k}})$ (2)
where the responses from $D^{L_{k}}$ pass the unit tests of correctness and instruction-following. We can directly optimize the LLM by generating answers and unit tests.
3 Experiments
3.1 Experiment Setup
Evaluation Metric
To check whether the model-generated response is correct (‘Corr.’: Correctness Pass@111) and follows the user instructions (‘Instr.’: Instruction-following Pass@111), we evaluate LLMs by executing the generated code with unit tests (pass rate for just one-time generation) and separately report ‘Corr.’ and ‘Instr.’ Pass@111 score in tables.
Code LLMs
We report 40+ popular LLMs for IFEvalCode, including o1-mini (o1_mini, ), GPT4o (gpt4, ), Claude-3.5/3.7 (claude, ; claude37, ), Deepseek-v3/R1 (deepseekv3, ; deepseek_r1, ) llama3 (llama3, ), qwen3 (qwen, ), and code-specific models like Qwen2.5-Coder (qwencoder, ), DeepSeekCoder (deepseek_coder, ), CodeLlama (code_llama, ), and Codegemma (codegemma, ). Additionally, we fine-tune Qwen2.5-Coder-7B as our baseline ControlledCoder.
ControlledCoder Training Setup
The training data for ControlledCoder comprises the constructed dataset IFEvalCode-Instruct and the Magicoder-Instruct code generation dataset (magicoder, ), ensuring fundamental instruction-following capabilities for code-related tasks. ControlledCoder, built on Qwen2.5-Coder-32B, is trained for 3 epochs using a cosine scheduler with an initial learning rate of $5\times 10^{-5}$ with a 3% warmup ratio. We employ AdamW (adamw, ) as the optimizer, with a batch size of 1024 and a maximum sequence length of 2048. (Details can be found in the Appendix).
Table 3: English evaluation of English queries in IFEvalCode of 8 programming languages. ‘Corr.’ denotes the correctness of the model-generated response, and ‘Instr.’ denotes the instruction-following accuracy of the response. The underlined values are the maximum in each column within each colored subgroup.
Models Params Python Java Cpp C-sharp Typescript Javascript Php Shell Avg.
Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr.
Closed-source LLMs
[…]
0.5B+ Open-source LLMs
Deepseek-Coder 1.3B 24.8 17.1 19.6 2.0 16.0 3.0 0.0 1.9 0.0 1.0 22.0 13.0 10.0 3.0 4.0 9.0 12.1 6.3
Qwen2.5-Coder 0.5B 18.1 19.0 8.8 2.0 8.0 2.0 9.7 1.9 2.0 18.0 13.0 12.0 11.0 4.0 3.0 9.0 9.3 8.5
Qwen2.5-Coder 1.5B 32.4 22.9 10.8 2.9 3.0 3.0 1.9 1.9 6.0 17.0 24.0 21.0 16.0 5.0 11.0 17.0 13.2 11.4
Qwen2.5-Coder 3B 35.2 24.8 11.8 2.0 11.0 5.0 11.7 1.9 0.0 18.0 26.0 29.0 22.0 12.0 15.0 23.0 16.7 14.4
Granite-Coder 3B 25.7 21.0 10.8 4.9 7.0 3.0 12.6 3.9 6.0 14.0 18.0 11.0 15.0 1.0 5.0 12.0 12.6 8.9
OpenCoder 1.5B 37.1 23.8 16.7 3.9 25.0 3.0 4.9 1.9 9.0 14.0 22.0 13.0 21.0 4.0 7.0 18.0 17.9 10.2
Yi-Coder 1.5B 32.4 21.0 13.7 2.0 13.0 4.0 11.7 2.9 8.0 17.0 26.0 21.0 16.0 4.0 8.0 14.0 16.2 10.7
Qwen3 0.6B 14.3 23.8 2.9 2.9 0.0 5.0 0.0 2.9 2.0 24.0 18.0 22.0 10.0 7.0 2.0 14.0 6.2 12.7
Qwen3-think 0.6B 13.3 15.2 2.0 2.0 1.0 5.0 0.0 1.9 2.0 6.0 8.0 3.0 13.0 0.0 0.0 6.0 4.9 4.9
Qwen3 1.7B 23.8 26.7 8.8 3.9 3.0 4.0 3.9 1.9 12.0 22.0 22.0 16.0 19.0 9.0 7.0 23.0 12.5 13.3
Qwen3-think 1.7B 25.7 21.0 3.9 2.9 3.0 2.0 0.0 2.9 10.0 11.0 22.0 17.0 28.0 2.0 7.0 13.0 12.5 9.0
Qwen3 4B 26.7 28.6 7.8 5.9 7.0 6.0 1.9 1.9 16.0 25.0 29.0 28.0 28.0 13.0 7.0 38.0 15.4 18
Qwen3-think 4B 34.3 21.9 12.7 4.9 6.0 4.0 1.0 3.9 12.0 12.0 29.0 20.0 25.0 2.0 13.0 17.0 16.7 10.7
6B+ Open-source LLMs
CodeLlama 7B 18.1 22.9 5.9 3.9 3.0 4.0 2.9 2.9 7.0 19.0 20.0 22.0 13.0 9.0 10.0 21.0 10.0 13.1
Llama3.1 8B 40.0 16.2 8.8 0.0 13.0 3.0 10.7 1.9 12.0 20.0 27.0 19.0 26.0 7.0 7.0 18.0 18.1 10.6
Deepseek-Coder 6.7B 37.1 21.9 27.5 7.8 20.0 3.0 4.9 1.9 16.0 17.0 30.0 20.0 30.0 8.0 12.0 19.0 22.2 12.3
Yi-Coder 9B 46.7 27.6 20.6 3.9 21.0 3.0 12.6 2.9 17.0 20.0 38.0 28.0 38.0 8.0 17.0 23.0 26.4 14.6
Granite-Coder 8B 35.2 19.0 11.8 3.9 11.0 3.0 3.9 1.9 13.0 15.0 32.0 15.0 21.0 1.0 11.0 8.0 17.4 8.4
OpenCoder 8B 38.1 24.8 21.6 5.9 24.0 4.0 5.8 1.9 16.0 17.0 42.0 19.0 26.0 9.0 14.0 15.0 23.5 12.1
CodeQwen1.5 7B 33.3 26.7 17.6 5.9 10.0 4.0 4.9 1.9 14.0 23.0 33.0 22.0 29.0 9.0 17.0 18.0 19.9 13.8
Qwen2.5-Coder 7B 41.9 27.6 25.5 3.9 16.0 3.0 11.7 2.9 19.0 22.0 36.0 30.0 43.0 11.0 20.0 28.0 26.7 16.0
Qwen3 8B 35.2 33.3 13.7 7.8 12.0 4.0 8.7 4.9 0.0 0.0 30.0 26.0 30.0 19.0 9.0 27.0 17.4 15.3
Qwen3-think 8B 35.2 31.4 15.7 9.8 6.0 4.0 1.9 5.8 16.0 14.0 30.0 24.0 28.0 2.0 22.0 14.0 19.4 13.2
14B+ Open-source LLMs
CodeLlama 13B 23.8 26.7 11.8 6.9 6.0 5.0 1.0 5.8 7.0 21.0 26.0 5.0 14.0 8.0 9.0 23.0 12.3 12.7
Qwen2.5-Coder 14B 48.6 27.6 27.5 7.8 22.0 2.0 18.4 2.9 20.0 25.0 38.0 31.0 34.0 12.0 20.0 35.0 28.6 17.9
Qwen3 14B 36.2 33.3 25.5 8.8 20.0 3.0 8.7 3.9 23.0 22.0 37.0 31.0 36.0 20.0 18.0 53.0 25.6 21.9
Qwen3-think 14B 51.4 31.4 22.5 15.7 16.0 5.0 1.9 4.9 30.0 24.0 48.0 33.0 43.0 4.0 24.0 17.0 29.6 16.9
Granite-Coder 20B 33.3 17.1 15.7 3.9 5.0 5.0 4.9 1.9 15.0 15.0 26.0 19.0 22.0 4.0 11.0 9.0 16.7 9.4
32B+ Open-source LLMs & MoE LLMs
Granite-Coder 34B 37.1 21.0 16.7 2.9 13.0 4.0 5.8 1.9 11.0 15.0 24.0 16.0 24.0 3.0 10.0 3.0 17.8 8.4
CodeLlama 34B 16.2 21.9 9.8 4.9 5.0 3.0 8.7 2.9 8.0 9.0 23.0 2.0 16.0 6.0 9.0 17.0 12.0 8.4
Deepseek-Coder 33B 41.9 22.9 33.3 4.9 22.0 4.0 7.8 1.9 24.0 20.0 33.0 18.0 37.0 7.0 19.0 20.0 27.3 12.3
Qwen2.5-Coder 32B 42.9 33.3 22.5 6.9 20.0 6.0 8.7 4.9 29.0 30.0 47.0 30.0 42.0 16.0 27.0 39.0 29.9 20.7
Llama3.1 70B 42.9 33.3 22.5 6.9 20.0 6.0 8.7 4.9 29.0 30.0 47.0 30.0 42.0 16.0 27.0 39.0 29.9 20.7
Qwen3 32B 49.5 33.3 25.5 8.8 17.0 7.0 10.7 1.9 20.0 29.0 38.0 34.0 41.0 28.0 22.0 52.0 28.0 24.2
Qwen3-think 32B 49.5 33.3 25.5 8.8 17.0 7.0 10.7 1.9 20.0 29.0 38.0 34.0 41.0 28.0 22.0 52.0 28.0 24.2
Qwen3 3/30B 54.3 30.5 30.4 11.8 31.0 4.0 1.0 4.9 31.0 26.0 48.0 28.0 46.0 2.0 36.0 21.0 34.7 16.0
Qwen3-think 3/30B 35.2 33.3 16.7 11.8 14.0 6.0 1.9 4.9 32.0 30.0 45.0 33.0 46.0 4.0 34.0 22.0 28.0 18.1
Qwen3 22B/235B 50.5 33.3 28.4 9.8 14.0 2.0 1.9 4.9 20.0 24.0 43.0 29.0 42.0 16.0 26.0 35.0 28.3 19.3
Qwen3-think 22B/235B 48.6 32.4 21.6 10.8 20.0 12.0 2.9 6.8 21.0 15.0 37.0 22.0 47.0 4.0 35.0 15.0 29.1 14.8
Deepseek-V3 37/637B 51.4 36.2 29.4 9.8 23.0 4.0 2.9 1.9 31.0 32.0 51.0 31.0 54.0 24.0 33.0 35.0 34.4 21.7
Deepseek-R1 37/637B 58.1 37.1 37.3 12.7 38.0 5.0 4.9 5.8 46.0 35.0 55.0 28.0 61.0 6.0 45.0 20.0 43.1 18.8
ControlledCoder 32B 58.3 38.2 39.3 13.5 30.0 12.0 10.7 6.8 44.0 38.0 47.0 38.0 48.0 30.0 45.0 52.0 40.3 34.2
Table 4: Chinese evaluation results in IFEvalCode of 8 programming languages. ‘Corr..’ denotes the correctness of the model-generated response, and ‘Instr.’ denotes the instruction-following accuracy of the response. The underlined fonts denote the best performance in the same parameter range.
Models Params Python Java Cpp C-sharp Typescript Javascript Php Shell Avg.
Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr.
Closed-source LLMs
[…]
0.5B+ Open-source LLMs
Deepseek-Coder 1.3B 22.9 10.5 14.7 1.0 7.0 2.0 1.9 1.9 0.0 1.0 16.0 13.0 7.0 7.0 1.0 2.0 8.9 4.8
Qwen2.5-Coder 0.5B 15.2 16.2 10.8 2.0 8.0 3.0 7.8 1.9 1.0 14.0 13.0 15.0 7.0 7.0 1.0 11.0 8.0 8.8
Qwen2.5-Coder 1.5B 28.6 21.9 9.8 2.9 2.0 5.0 0.0 1.9 6.0 20.0 28.0 23.0 18.0 11.0 4.0 24.0 12.1 13.7
Qwen2.5-Coder 3B 31.4 20.0 13.7 3.9 12.0 5.0 11.7 2.9 7.0 19.0 31.0 23.0 27.0 12.0 10.0 30.0 18.0 14.4
Granite-Coder 3B 25.7 21.0 10.8 4.9 7.0 3.0 12.6 3.9 6.0 14.0 18.0 11.0 15.0 1.0 5.0 12.0 12.6 8.9
OpenCoder 1.5B 39.0 16.2 21.6 3.9 20.0 2.0 6.8 1.9 10.0 14.0 18.0 14.0 21.0 8.0 7.0 17.0 18.0 9.6
Yi-Coder 1.5B 34.3 17.1 15.7 1.0 16.0 5.0 14.6 1.9 10.0 11.0 22.0 21.0 15.0 7.0 2.0 13.0 16.3 9.6
Qwen3 0.6B 11.4 21.0 2.9 3.9 1.0 6.0 0.0 1.9 0.0 25.0 13.0 19.0 8.0 15.0 0.0 22.0 4.6 14.2
Qwen3-think 0.6B 10.5 10.5 2.9 3.9 0.0 2.0 0.0 1.9 1.0 10.0 9.0 8.0 6.0 1.0 0.0 7.0 3.7 5.6
Qwen3 1.7B 27.6 29.5 6.9 3.9 4.0 5.0 1.0 1.9 9.0 25.0 19.0 19.0 15.0 13.0 5.0 31.0 11.0 16.0
Qwen3-think 1.7B 21.0 14.3 2.9 2.9 0.0 0.0 0.0 1.0 10.0 1.0 14.0 3.0 21.0 2.0 3.0 3.0 9.0 3.5
Qwen3 4B 24.8 30.5 6.9 7.8 5.0 4.0 2.9 3.9 14.0 26.0 28.0 31.0 29.0 20.0 8.0 47.0 14.8 21.2
Qwen3-think 4B 24.8 7.6 4.9 1.0 1.0 1.0 0.0 0.0 9.0 5.0 16.0 3.0 21.0 1.0 14.0 6.0 11.4 3.1
6B+ Open-source LLMs
CodeLlama 7B 0.0 1.9 5.9 1.0 1.0 3.0 0.0 2.9 5.0 2.0 23.0 0.0 1.0 2.0 2.0 17.0 4.7 3.7
Llama3.1 8B 35.2 21.9 9.8 2.0 13.0 2.0 15.5 1.9 10.0 19.0 24.0 23.0 24.0 14.0 5.0 24.0 17.2 13.5
Deepseek-Coder 6.7B 38.1 23.8 20.6 3.9 16.0 3.0 2.9 1.9 18.0 16.0 34.0 17.0 25.0 10.0 16.0 23.0 21.4 12.3
Yi-Coder 9B 43.8 24.8 31.4 5.9 21.0 2.0 7.8 1.9 18.0 16.0 33.0 24.0 46.0 11.0 14.0 23.0 26.9 13.6
Granite-Coder 8B 32.4 19.0 11.8 3.9 14.0 4.0 6.8 1.9 9.0 13.0 29.0 19.0 9.0 3.0 10.0 16.0 15.3 10.0
OpenCoder 8B 40.0 25.7 17.6 4.9 16.0 5.0 4.9 1.9 15.0 19.0 30.0 19.0 32.0 12.0 13.0 26.0 21.1 14.2
CodeQwen1.5 7B 41.0 21.9 15.7 4.9 12.0 4.0 9.7 1.9 12.0 23.0 33.0 25.0 27.0 11.0 13.0 23.0 20.5 14.3
Qwen2.5-Coder 7B 42.9 27.6 21.6 4.9 11.0 5.0 6.8 2.9 16.0 21.0 39.0 29.0 36.0 11.0 21.0 32.0 24.3 16.7
Qwen3 8B 36.2 32.4 11.8 5.9 7.0 5.0 5.8 3.9 0.0 0.0 29.0 32.0 32.0 27.0 6.0 18.0 16.0 15.6
Qwen3-think 8B 28.6 12.4 8.8 2.9 3.0 1.0 1.9 1.0 14.0 2.0 26.0 6.0 27.0 0.0 15.0 6.0 15.6 4.0
14B+ Open-source LLMs
CodeLlama 13B 19.0 22.9 10.8 3.9 7.0 4.0 4.9 2.9 7.0 24.0 24.0 8.0 22.0 7.0 4.0 22.0 12.3 11.9
Qwen2.5-Coder 14B 47.6 29.5 26.5 8.8 22.0 3.0 19.4 2.9 24.0 32.0 41.0 30.0 39.0 15.0 19.0 36.0 29.9 19.6
Qwen3 14B 39.0 32.4 19.6 8.8 17.0 3.0 10.7 4.9 23.0 31.0 41.0 33.0 40.0 23.0 18.0 53.0 26.0 23.6
Qwen3-think 14B 49.5 26.7 16.7 2.9 10.0 1.0 2.9 1.0 23.0 8.0 34.0 6.0 37.0 1.0 22.0 10.0 24.4 7.2
Granite-Coder 20B 30.5 21.9 15.7 2.9 4.0 4.0 6.8 3.9 13.0 13.0 25.0 19.0 26.0 3.0 4.0 10.0 15.7 9.8
32B+ Open-source LLMs & MoE LLMs
Granite-Coder 34B 34.3 23.8 12.7 4.9 10.0 4.0 5.8 1.9 16.0 15.0 29.0 23.0 24.0 4.0 5.0 9.0 17.2 10.7
CodeLlama 34B 4.8 15.2 5.9 2.0 15.0 3.0 12.6 1.9 10.0 6.0 18.0 0.0 17.0 10.0 3.0 22.0 10.7 7.5
Deepseek-Coder 33B 44.8 21.0 26.5 4.9 24.0 7.0 7.8 1.9 21.0 18.0 35.0 20.0 30.0 14.0 20.0 26.0 26.2 14.1
Qwen2.5-Coder 32B 43.8 34.3 28.4 8.8 19.0 2.0 12.6 5.8 25.0 29.0 41.0 34.0 44.0 24.0 23.0 38.0 29.6 22.0
Llama3.1 70B 47.6 22.9 14.7 2.0 19.0 3.0 11.7 2.9 19.0 24.0 35.0 28.0 39.0 19.0 21.0 40.0 25.9 17.7
Qwen3 32B 42.9 41.0 24.5 7.8 17.0 7.0 9.7 3.9 27.0 25.0 38.0 31.0 40.0 23.0 19.0 51.0 27.3 23.7
Qwen3-think 32B 54.3 19.0 30.4 5.9 24.0 2.0 7.8 1.9 26.0 2.0 42.0 7.0 50.0 1.0 30.0 6.0 33.1 5.7
Qwen3 3/30B 43.8 33.3 28.4 11.8 19.0 6.0 5.8 5.8 24.0 31.0 42.0 31.0 38.0 29.0 21.0 48.0 27.8 24.4
Qwen3-think 3/30B 42.9 21.0 16.7 2.9 12.0 0.0 2.9 2.9 21.0 4.0 41.0 10.0 40.0 1.0 22.0 7.0 24.8 6.2
Qwen3 22B/235B 46.7 38.1 22.5 8.8 12.0 3.0 1.0 2.9 25.0 24.0 44.0 30.0 47.0 22.0 20.0 38.0 27.3 20.9
Qwen3-think 22B/235B 46.7 11.4 26.5 2.0 26.0 1.0 1.0 0.0 32.0 1.0 43.0 2.0 43.0 1.0 36.0 2.0 31.7 2.6
Deepseek-V3 37/671B 50.5 35.2 32.4 9.8 21.0 5.0 4.9 1.9 32.0 34.0 51.0 25.0 51.0 29.0 34.0 38.0 34.6 22.2
Deepseek-R1 37/671B 58.1 37.1 37.3 12.7 38.0 5.0 4.9 5.8 46.0 35.0 55.0 28.0 61.0 6.0 45.0 20.0 43.1 18.8
ControlledCoder 32B 54.8 39.2 40.3 13.8 34.0 7.0 11.7 6.8 48.0 35.0 49.0 40.0 48.0 32.0 51.0 55.0 42.1 28.6
3.2 Main Results
Results of English queries in IFEvalCode
Table 3 presents evaluation results of “Correctness” (Corr.) and “Instruction Compliance” (Instr.) for 40+ LLMs across 888 programming languages (Python, Java, C++, C#, TypeScript, JavaScript, PHP, Shell). These models vary in size and origin (closedsource/opensource/mixed expert models) and are tested on English programming queries. Closed-source LLMs (such as Claude-3.7, GPT-4.1, and various mini versions) generally achieve overall correctness rates concentrated between 30% and 40%. Instruction compliance rates typically hover around 20% to 25%. Open-source LLMs with 0.5B to 3B parameters generally demonstrate limited performance, with average correctness rates mostly between 6% and 18%. The top performer in this range, OpenCoder (1.5B), only reaches 17.9%. This indicates that small-scale models still struggle to achieve satisfactory results in cross-language, multi-task coding scenarios. Open-source models in the 6B to 14B parameter range show significant improvement. The 7B Qwen2.5-Coder, 8B Qwen3-think, and 14B Qwen3 models all surpass average correctness rates of 25% to 30%. Notably, the 14B Qwen3-think achieves a multilingual overall correctness rate of 29.6%. Among models with 32B parameters or more / MoE architectures, Deepseek-R1 (37/637B) performs the best, reaching an average correctness rate as high as 43.1%. It sets the highest group scores across multiple languages: Python (58.1%), Java (37.3%), C++ (38.0%), TypeScript (46.0%), JavaScript (55.0%), PHP (61.0%), and Shell (45.0%). Its instruction compliance rate also improves to around 19% to 20%. Overall, model scale and architectural complexity (such as MoE expert models) significantly boost programming task capabilities. At the same time, whether closed-source or open-source, the current instruction compliance rates remain noticeably lower than correctness rates. This highlights the need for further optimization in instruction tuning and code execution verification in the future.
Results of English queries in IFEvalCode
Table 4 presents the evaluation results of various large models on Chinese programming queries (covering eight languages: Python, Java, C++, C#, TypeScript, JavaScript, PHP, and Shell) in terms of "Correctness" (Corr.) and "Instruction Compliance" (Instr.). The best performers within each size range are highlighted in bold. Closed-source models (Claude-3.x, GPT-4.0/4.1, and their mini versions) have an overall average correctness between 27% and 38%, with instruction compliance around 20% to 25%. Among them, GPT-4.1 leads the closed-source group in both correctness (36.4%) and instruction compliance (24.7%). Although o1-mini and o3-mini show higher correctness rates (39.3% and 37.8%, respectively), their instruction compliance rates are relatively lower (23.5% and 24.1%). Open-source models in the 0.5B–3B parameter range show limited performance, with average correctness mostly between 5% and 18%. OpenCoder (1.5B) at 18.0% and Qwen2.5-Coder (3B) also at 18.0% lead this group, indicating that small-scale models still need improvement in multilingual coding tasks. In the 6B–14B parameter range, open-source models begin to show significant improvement—Yi-Coder (9B) achieves an average correctness of 26.9%, while Qwen2.5-Coder (14B) ranks first in this group with a correctness of 29.9% and instruction compliance of 19.6%. Qwen3 (8B) also surpasses a 23% correctness rate and obtains 32% instruction compliance on Python and TypeScript. Among large models with 32B+ parameters and MoE (Mixture of Experts) architectures, Deepseek-R1 (37/671B) demonstrates the most outstanding performance, with an average correctness of 43.1% and the highest scores across multiple sub-tasks including Python, Java, C++, TypeScript, JavaScript, PHP, and Shell. Although its instruction compliance (18.8%) is not the highest, it remains relatively high. Deepseek-V3 follows closely (34.6%), with Qwen2.5-Coder 32B (29.6%) and Llama3.1-70B (25.9%) ranking next. Overall, model scale and architecture (especially MoE expert models) have a significant positive impact on Chinese programming capabilities. Meanwhile, instruction compliance rates are generally lower than correctness rates across all models, highlighting the need for further optimization in instruction tuning and multilingual semantic understanding in the future.
4 Further Analysis
Ablation Study.
Table 5: Ablation study.
ID Method Corr.avg Instr.avg
Qwen2.5-Coder-32B 29.9 20.7
+ $D_{s_{1}}$ 34.5 26.4
+ $D_{s_{2}}$ 38.2 30.4
+ Iter2 40.3 34.2
+ Iter3 40.5 33.2
In Table 5, we conduct the ablation study to explore the effect of each component in IFEvalCode-Instruct. ControlledCoder is fine-tuned on the base Qwen2.5-Coder by combining the instruction corpus from the forward and backward constraints generation ( and outperforms the baseline Qwen2.5-Coder ). The fine-tuned coder ControlledCoder can iteratively update the instruction corpus IFEvalCode-Instruct using forward and backward constraints generation. We can see that joint multilingual rejecting fine-tuning can effectively bring improvement for controlled code generation.
(a) Distribution of constraints number
(b) Scores of different constraints number
Figure 5: Figure 5(a) and 5(b) plot the trends with the number of constraints increaing
Correlation between Correctness and Code Instruction Following
Figure 6 displays the correctness overlap between correctness score (Eh), instruction following score (Zh), correctness score (En), and instruction following score (Zn) of 888 different programming languages.
Figure 6: The Venn diagram of the numbers of the correct samples for each programming language.
Figure 6 shows a grid of four-set Venn diagrams—one subplot per programming language—where each diagram partitions the model’s outputs into the following four evaluation sets: (1) Chinese-prompt correctness, (2) Chinese-prompt instruction compliance, (3) English-prompt correctness, and (4) English-prompt instruction compliance. By inspecting the overlaps, we can immediately see several consistent patterns across all eight languages: Correctness dominates. Both the Chinese-correctness and English-correctness sets are substantially larger than their corresponding instruction-compliance sets, indicating that LLMs more often produce functionally correct code than code that also respects the extra constraints. Cross-language consistency in correctness. The overlap between Chinese-correctness and English-correctness is very large in every language, showing that whether the prompt is given in Chinese or English has little effect on a model’s ability to generate a passing solution. Instruction compliance is more fragile. The Chinese-instruction and English-instruction sets are much smaller, and their overlap is modest. In other words, even if a model is correct in both languages, it frequently fails to satisfy the same set of style or structural constraints when the prompt language switches. The smallest region is the four-way intersection. Only a small fraction of examples are simultaneously correct and compliant in both Chinese and English, demonstrating that achieving full controlled generation across languages remains the hardest challenge.
5 Related Work
Code Large Language Model
The rapid progress of large language models(gpt4, ; llama2, ; llama3, ; deepseekv3, ; deepseek_r1, ) has enabled complex code-related tasks. Early models (e.g. BERT (bert, ) and GPT (gpt, )) trained on billions of code snippets, focused on code understanding and generation (codex, ; code_bert, ; bloom, ; AlphaCode, ; codet5, ; santacoder, ). Recent advances in domain-specific pre-training and instruction fine-tuning (kun, ; mammoth2, ) have enhanced models like Qwen2.5-Coder (qwen_coder, ) and Deepseek-Coder (deepseek_coder, ), achieving strong performance in code completion, synthesis, and repair.
Code Benchmarks
In the field of code, a rich tapestry of benchmarks (codex, ; mbpp, ; humaneval_x, ; codereval, ; arcade_nl2code, ; humaneval_xl, ; xcodeeval, ; BabelCode, ) has been developed to address the challenges of assessing code quality, functionality, and efficiency. Current benchmarks support the evaluation of various tasks, including code generation (e.g. NCB (naturalcodebench, ) and MultiPL-E (multipl_e, )), code repair  (e.g. DebugBench (debugbench, ) and MdEval (mdeval, )), code competition (e.g. LCB (livecodebench, ) and CodeElo (codeelo, )), software development (e.g. Aider and SWE-Bench (swe_bench, )). But existing code benchmarks ignores the multilingual evaluation of controlled code generation for LLMs. IFEval (ifeval, ) emphasizes the importance of evaluating the proficiency of LLMs in instruction following, where IFEval contains a series of “verifiable instructions”.
6 Conclusion
In this paper, we first introduce IFEvalCode, a multilingual benchmark with 1.6K samples across 8 programming languages, designed to evaluate both code correctness and code controllability in LLM-generated code. To address these challenges, we propose a novel framework that combines forward constraints generation and backward generation, enhancing LLMs’ ability to follow human instructions for controlled code generation. Our approach advances the reliability and precision of LLMs in adhering to user intent while producing functionally accurate code.
Appendix A List of Large Language Models
Table 6 lists all the large language models evaluated in this study and their download links; we then systematically evaluate the performance of these models on the IFEvalCode benchmark.
Table 6: Evaluation Model List
Model Download Link
Claude-3.7-Sonnet https://www.anthropic.com/news/claude-3-7-sonnet
Claude-3.5-Sonnet https://www.anthropic.com/news/claude-3-5-sonnet
GPT-4o https://platform.openai.com/docs/models/gpt-4o
GPT-4o-mini https://platform.openai.com/docs/models/gpt-4o-mini
GPT-4.1 https://platform.openai.com/docs/models/gpt-4.1
GPT-4.1-mini https://platform.openai.com/docs/models/gpt-4.1-mini
o1-mini https://platform.openai.com/docs/models/o1-mini
o3-mini https://platform.openai.com/docs/models/o3-mini
o4-mini https://platform.openai.com/docs/models/o4-mini
grok-3 https://docs.x.ai/docs/models#models-and-pricing
grok-3-mini-fast https://docs.x.ai/docs/models#models-and-pricing
Deepseek-Coder-1.3B https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-instruct
Qwen2.5-Coder-0.5B https://huggingface.co/Qwen/Qwen2.5-Coder-0.5B
Qwen2.5-Coder-1.5B https://huggingface.co/Qwen/Qwen2.5-Coder-1.5B
Qwen2.5-Coder-3B https://huggingface.co/Qwen/Qwen2.5-Coder-3B
Granite-Coder-3B https://huggingface.co/ibm-granite/granite-3b-code-base-128k
OpenCoder-1.5B https://huggingface.co/infly/OpenCoder-1.5B-Instruct
Qwen3-0.6B https://huggingface.co/Qwen/Qwen3-0.6B
Qwen3-1.7B https://huggingface.co/Qwen/Qwen3-1.7B
Qwen3-4B https://huggingface.co/Qwen/Qwen3-4B
CodeLlama-7B https://huggingface.co/meta-llama/CodeLlama-7b-Instruct-hf
Deepseek-Coder-6.7B https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct
Yi-Coder-9B https://huggingface.co/01-ai/Yi-Coder-9B-Chat
Granite-Coder-8B https://huggingface.co/ibm-granite/granite-8b-code-instruct-128k
OpenCoder-8B https://huggingface.co/infly/OpenCoder-8B-Instruct
CoderQwen1.5-7B https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat
Qwen2.5-Coder-7B https://huggingface.co/Qwen/Qwen2.5-Coder-7B
Qwen3-8B https://huggingface.co/Qwen/Qwen3-8B
Qwen2.5-Coder-14B https://huggingface.co/Qwen/Qwen2.5-Coder-14B
Qwen3-14B https://huggingface.co/Qwen/Qwen3-14B
Granite-Coder-20B https://huggingface.co/ibm-granite/granite-20b-code-instruct-8k
Granite-Coder-34B https://huggingface.co/ibm-granite/granite-34b-code-instruct-8k
Deepseek-Coder-33B https://huggingface.co/deepseek-ai/deepseek-coder-33b-instruct
Qwen2.5-Coder-32B https://huggingface.co/Qwen/Qwen2.5-Coder-32B
Qwen3-32B https://huggingface.co/Qwen/Qwen3-32B
Qwen3-30B-A3B https://huggingface.co/Qwen/Qwen3-30B-A3B
Qwen3-235B-A22B https://huggingface.co/Qwen/Qwen3-235B-A22B
DeepSeek-V3 https://huggingface.co/deepseek-ai/DeepSeek-V3
DeepSeek-R1 https://huggingface.co/deepseek-ai/DeepSeek-R1
Appendix B Limitations
Programming Languages
To ensure a more comprehensive evaluation of multilingual controlled code generation, it would be beneficial to expand IFEvalCode to include additional programming languages, particularly those that are less commonly used or serve niche applications, as the current coverage of eight languages may not fully represent the diversity and specialized use cases present in real-world software development.
Complex Controlled Code Generation
While IFEvalCode aims to simulate realistic controlled code generation scenarios, IFEvalCode ignores the complex controlled code generation and lacks repository-based evaluation.
Appendix C Ethical Considerations
C.1 Potential Risks
IFEvalCode serves as a comprehensive assessment tool for controlled code generation. However, improper use may lead to risks such as incorrect program analysis, faulty repairs, or even system failures. To mitigate these risks, we strongly recommend conducting evaluations within a sandbox environment, which effectively isolates potential hazards and safeguards both the accuracy of the assessment and the stability of the underlying system.
Appendix D Human Annotation
Annotators
To construct IFEvalCode, we developed a rigorous and systematic human annotation process, guided by meticulously designed annotation protocols, to ensure the accuracy, consistency, and high quality of questions. This process integrates multiple quality control measures at every stage to maintain reliability and precision across all collected data. We recruited 666 programmers with computer science degrees equipped with strong computer science fundamentals.
Appendix E Related Work
Code Large Language Models.
Large language models (LLMs) designed for coding tasks have demonstrated exceptional capabilities in code generation, debugging, translation, and other essential functions for modern software engineering [14, 4, 43, 57], with numerous benchmarks developed to evaluate these capabilities, though many focus on a limited selection of programming languages like Python and Java [83, 6, 28]. Recent advancements in code LLMs, such as Code Llama [53], DeepSeek-Coder [22], OpenCoder [23], and Qwen2.5-Coder [26], have significantly improved multilingual code generation and debugging, evaluated using benchmarks like MultiPL-E [10], McEval [11], and MdEval [39]. Code generation, a fundamental task for code LLMs, requires interpreting natural language descriptions and generating corresponding code snippets [20, 31, 37, 34], with benchmarks proposed for code translation [69], code completion [8, 78], and structured data understanding [66, 56].
Instruction Following.
Automatic program debugging holds substantial practical value. With the emergence of LLM capabilities, a growing number of individuals are utilizing LLMs for code debugging, leading to extensive research in this field. Code Debugging includes serval tasks such as bug or vulnerability detection [47, 3, 75, 77, 85], fuzz test [15, 67, 71], program repair  [65, 35, 79, 49, 19, 58, 63], GitHub issues auto resolving  [29, 12, 59]. To effectively assess the code debugging capabilities of LLMs, several benchmark tests have been introduced  [48, 55, 68, 79, 60, 72]. Notably, DebugBench [60] provides a comprehensive classification of error types and analyzes the debugging capabilities of LLMs based on these categories. Similarly, DebugEval [72] has designed various debugging-related tasks to evaluate LLM performance across different task dimensions. However, these studies focus on 1 to 3 languages. In reality, there are significant differences in code errors between languages, leading to numerous language-specific errors. To address this gap, we propose IFEvalCode, a comprehensive code debugging benchmark covering 20 languages, aiming to assess LLM debugging capabilities from a broader perspective.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Please see the abstract and introduction section.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please check Limitations section in Appendix B.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please check Experiments section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will open source our training code and training data soon, and have provided the model training code in the supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please check Experiments section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Experimental statistical significance is not applicable to our experiments. In section 3.1, we report the evaluation metrics and method.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please check section 3.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we have read the Code of Ethics in detail.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please check Appendix C.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: For the data sources of synthetic data, we have filtered out the risk information.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We comply with the license for the data and models used in training and evaluation.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes, we provide training instructions in the code of the supplementary material. We will also add detailed instructions when we open source our data.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Our benchmark curation involves some text data annotation work, and we provide compensation in accordance with regulations. Please check Appendix. The data does not involve risky content or illegal information. The main content of the data is common programming questions.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: Yes. Our work does not involve human experiments. The text data annotation work involved complies with relevant regulations. The data does not involve risky content or illegal information. The main content of the data is common programming questions.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We describe the usage of LLM in the experimental setup section.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.