### v2

# IFEvalCode: Controlled Code Generation

Abstract
Code large language models (Code LLMs) have achieved significant advancements in various code-related tasks, particularly in code generation, where the code LLMs produce the target code from natural language descriptions. However, in realistic scenarios, users often expect the returned code to strictly follow the given detailed requirements in many aspects (e.g. the style of code, the number of code lines, or the number of lines), instead of only requiring the correctness of the generated code. Controlled code generation means that the generated response from code LLMs should adhere to specific human guidelines or standards, whereas the LLM should have a strong instruction-following capability in the field of the code. In this paper, we propose forward constraints generation and backward constraints generation for controlled code generation to enhance the capability of LLM in following human instructions. Then, we build a multilingual benchmark IFEvalCode to evaluate the code instruction-following capability of the LLMs. IFEvalCode consists of 1.6K samples (Python, Java, Javascript, Typescript, Shell, Cpp, Php, and C-sharp) and each test sample contains the Chinese and English query. Different from the existing code benchmarks, we separately design the test function to verify the correctness of the code (Corr.) and whether the generated code follows the human instruction (Instr.). Extensive experimental results on IFEvalCode of 40+ LLMs emphasize that closed-source LLMs still dominate in controllable code generation compared to open-source LLMs, and the ability of LLMs to generate controllable code is far behind its ability to generate correct code.
1 Introduction
Large Language Models (LLMs) have made significant strides in code-related tasks, particularly in producing target code snippets from the given human description. It has captured considerable attention from academia and industry, owing to its practical relevance in software development. The current top-tier closed-source LLMs, such as GPT4.5 (gpt45, ) and Claude3.7 (claude37, ), can correctly generate the file-level code snippet most of the time. Besides, a series of open-source code LLMs, such as StarCoder1/2 (starcoder, ; starcoder2, ), Deepseek-Coder (deepseek_coder, ), OpenCoder (opencoder, ), and Qwen2.5-Coder (qwen_coder, ), has also become a force that cannot be ignored by industry and academia of code intelligence.
Most existing code benchmarks (mceval, ), such as HumanEval (codex, ), BigCodeBench (bigcodebench, ), and MBPP (mbpp, ) are designed to evaluate the correctness of generated code by code execution with the unit tests. Further, some benchmarks, such as Chatbot Arena (llm_as_a_judge, ) and CodeArena (codearena, ) are proposed to evaluate the alignment between the model-generated response and human preference using the LLM-as-a-Judge. For the general domain, IFEval (ifeval, ) is proposed to evaluate the proficiency of LLM in controlled text generation, where the instructions are amenable to objective verification of compliance. In the field of code, in addition to requiring the generation of correct code, users often also require the generation of code meeting various objective requirements, such as code style, variable naming, specific algorithms, time complexity, etc. Therefore, we try to explore the proficiency of LLMs in controlled code generation by designing a framework to enhance and evaluate the capabilities in code instruction following.
Figure 1: Comparision between general code generation and Controllable code generation for 8 programming language and 2 natural language.
To evaluate the capability of generating the controlled code, we construct a multilingual benchmark using human annotators, IFEvalCode, comprising nearly 1.6K samples across 888 programming languages (Python, Java, CPP, C-Sharp, TypeScript, JavaScript, PHP, Shell), with each sample including both Chinese and English queries. Unlike existing benchmarks, IFEvalCode uniquely assesses two critical aspects: the correctness (code correctness) of the generated code and its adherence to human instructions (code controllability). In this paper, we introduce a novel framework combining forward constraints generation and backward generation for controlled code generation, aimed at enhancing the ability of LLM to follow human instructions in controlled code generation.
The contributions are summarized as follows:
  • We introduce IFEvalCode, a multilingual benchmark comprising 1,620 samples across 888 programming languages (Python, Java, JavaScript, TypeScript, Shell, C++, and C#) with each sample featuring paired Chinese and English queries to facilitate comprehensive evaluation in controlled code generation.
  • We integrate forward constraints generation and backward constraints generation to generate the instruction corpus IFEvalCode-Instruct, designed to enhance the capability of LLMs in adhering to human instructions during code generation. By jointly training to improve the controllability, accuracy, and alignment of generated code with specific user requirements, thereby advancing the performance of LLMs in structured and instruction-driven coding tasks.
  • We perform a comprehensive evaluation and analysis of existing 40+ LLMs on IFEvalCode, and several insightful findings are as follows: (1) Controlled code generation is challenging for LLMs. LLMs still have huge room for improvement and a long journey ahead to reach their full potential. (2) Closed-source models (e.g. Claude3.7) and LLMs with large sizes still significantly outperform open-source LLMs and LLMs with small sizes in controllable code generation. (3) The ability to generate specified code is significantly weaker than the ability to generate correct code.
2 Multilingual Controlled Code Generation
2.1 Task Definition
Given the $k$-th programming language $L_{k}\in\{L_{k}\}_{k=1}^{K}$ (in this work, $K=8$). We provide the problem description $q^{L_{k}}$ and instruction constraints $c^{L_{k}}=\{c^{L_{k}}_{1},\dots,c^{L_{k}}_{m}\}$ as input to the code LLMs $\mathcal{M}$ (Note: the constraints for code generation are language-specific). The LLMs generate the corresponding code $a^{L_{k}}$ by sampling from the code generation distribution $P(a^{L_{k}}|q^{L_{k}},c^{L_{k}};\mathcal{M})$. The generated code is then executed with the provided test cases $u_{L^{k}}$ to verify correctness. Specifically, we design another check function $v^{L_{k}}$ for each query to determine whether the generated code is controllable, meeting human instruction constraints. This process ensures that the generated code $a^{L_{k}}$ functionally aligns with the problem requirements and :
$r^{L_{k}}_{u}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}};\mathcal{M});u^{L_{k}});\ \ \ \ \ \ r^{L_{k}}_{v}=\mathbb{I}(P(a^{L_{k}}|q^{L_{k}};\mathcal{M});v^{L_{k}});$ (1)
where $\mathbb{I}(\cdot)$ is the code execution in a sandbox environment with the given test cases $u^{L_{k}}$ or the check function $v^{L_{k}}$. when the generated code $a^{L_{k}}$ passes the unit tests $u^{L_{k}}$, the evaluation result $r^{L_{k}}_{u}=1$, else $r^{L_{k}}_{u}=0$. For controlled code generation, the check function $v^{L_{k}}$ is used to check whether the generated code satisfies the human instruction. $u^{L_{k}}$ and $v^{L_{k}}$ separately denote the scores of code correctness and code controllability.
2.2 Overview of IFEvalCode
Table 1: Benchmark statistics.
Statistics Number (Zh/En)
Problems $/810$ 810
Python $/103$ 103
Java $/102$ 102
Cpp $/100$ 100
Typescript $/100$ 100
Javascript $/100$ 100
Php $/100$ 100
Shell $/100$ 100
C-Sharp $/100$ 100
#Instruction Constraints
- min/max/avg length 2/15/5
Length
Chinese Question
min/max/avg 142142142/1,3731,3731 , 373/358358358 tokens
English Question
min/max/avg 135135135/1,3041,3041 , 304/340340340 tokens
In Table 1, IFEvalCode consists of 1.6K unique problems for 8 programming languages. The benchmarks can be used for bilingual natural languages, including Chinese and English (1.6K for English and 1.6K translated queries for Chinese). Each sample in IFEvalCode includes (en_question, zh_question, check_instruction, check_correctness), where ‘check_instruction‘ is used to check the code correctness and ‘check_correctness‘ judges whether the generated code follows the human instruction. We calculate the length of the question using the Qwen2.5-Coder tokenizer (qwen_coder, ) and count the number of instruction constraints for each sample. Each question contains 3 constraints and 100 tokens on average, where each prompt contains a problem description and the corresponding function or class signature.
In Table 2, we compare IFEvalCode with other code evaluation benchmarks in 6 aspects. Considering the evaluation cost and effectiveness, IFEvalCode has 1620 samples for 8 programming languages and 2 human languages. Specifically, we introduce the two parts of unit tests to check the code correctness and whether the code follows the human instruction.
Table 2: Comparison between IFEvalCode and other code benchmarks. Different from the existing benchmarks, IFEvalCode effectively provides a new comprehensive view of controlled code generation by adding the ‘check_instruction‘ part into the test samples.
Benchmark #Languages #Domains Size (Zh/En) Difficulty Human Check Check Code Controllability
HumanEval (codex, ) Python 3 164 (En) [Uncaptioned image]
MBPP (mbpp, ) Python Algorithm 974 (En) [Uncaptioned image]
BigCodeBench (bigcodebench, ) Python 7 2000+ (En) [Uncaptioned image] [Uncaptioned image]
CodeArena (codearena, ) 44 7 397+ (En) [Uncaptioned image] [Uncaptioned image]
FullStackBench (fullstackbench, ) 16 11 3374 (En) [Uncaptioned image] [Uncaptioned image]
LiveCodeBench (livecodebench, ) Python 3 450+ (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MultiPL-E (multipl_e, ) 19 3 12,667 (En) [Uncaptioned image] [Uncaptioned image]
McEval (mceval, ) 20 10+ 16K (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MdEval (mdeval, ) 40 10+ 39K (En) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DS-1000 (ds1000, ) Python Data Size 1000 (En) [Uncaptioned image] [Uncaptioned image]
IFEvalCode (Ours) 8 3 810/810 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(a) Domain types.
(b) Types of instruction constraints.
Figure 2: Figure 2(a) shows domain types and Figure 2(b) lists instruction types.
2.3 IFEvalCode Construction & Quality Control
To create IFEvalCode for controlled code generation, we propose a systematic human annotation and check procedure, which is guided by defined guidelines to guarantee accuracy and consistency. Figure 4 illustrates the overall process of dataset construction. We begin by collecting code snippets and code-related documents from websites and recruit 6 computer science graduates as annotators. The annotators must follow the provided guidelines (1) ensure the diversity of the questions. (2) the question and instruction constraints should be challenging for existing LLMs. (3) translate the English question into the Chinese question. To increase the difficulty of IFEvalCode, the annotator filter out the questions, which half LLMs (GPT4o, DeepSeek-V3, Claude3.7, etc) can correctly answer.
Figure 3: Examples of the verifiable instructions with ‘check_instruction’ in IFEvalCode.
Figure 4: Construction of the benchmark IFEvalCode and instruction corpus IFEvalCode-Instruct.
2.4 Reject-sampling Fine-tuning for Controlled Code Generation
Creating Queries and Unit Tests with Constraints
To create the instruction corpora, we collect the code-related websites and remove the HTML tags to obtain the code snippet with the corresponding natural language description as the seed data. Then, we select the select seed data from 8 languages and force the code LLM (Qwen2.5-Coder 32B) to generate a challenging question $q^{L_{k}}$ of programming language $L_{k}$. To verify the correctness of the response, we prompt the LLM to generate the test cases $u_{k}$ to create ‘check_correctness’. To independently check whether the model-generated response satisfies the constraints in the question, we assume that the response is correct and only check whether the generation code follows the user instructions. Therefore, for each query $q^{L_{k}}$ with constraints $c^{L_{k}}$, we create the unit tests $r_{u}^{L_{k}}$ to check the correctness of the response and $r_{u}^{L_{k}}$ to check the whether the generated code is under control.
Forward Constraints Generation
Given the recalled code-related documents from Common Crawl, we adopt Qwen2.5-Coder-32B to create new questions by drawing inspiration from the code-related documents for a general question. To effectively inject the additional instruction constraints into the question, we select the constraint types from the constraints types pool $C=\{c_{i=1}\}_{i=1}^{m}$ and force the model to generate a new question $q^{L_{k}}$ with constraints. After generating the question $q^{L_{k}}$ with constraints $c^{L_{k}}$, we force the LLM to generate $r^{L_{k}}_{u}$ and $r^{L_{k}}_{v}$ and the only keep the samples passing both unit tests. The generated corpus is denoted as $D_{s_{1}}$
Backward Constraints Generation
Since the forward constraints generation must first pre-define the constraint types and then generate the corresponding question, we introduce the backward constraints generation to synthesize the question with random instruction constraints. First, we prompt the LLM to generate the question $q^{L_{k}}$ and answer $a^{L_{k}}$. Then, we prompt the LLM to summarize which constraints the generated code complies with and obtain othe constraints $c^{L_{k}}$. Finally, we generate the unit tests $r^{L_{k}}_{u}$ and $r^{L_{k}}_{v}$, which are further used to verify the response and the created instruction corpus from backward constraints generation is denoted as $D_{s_{2}}^{L_{k}}$.
Joint Multilingual Rejecting Fine-tuning for Controlled Code Generation
Given the large-scale instruction corpus $D^{L_{k}}=D_{s_{1}}^{L_{k}}\bigcup D_{s_{2}}^{L_{k}}$ of different programming languages, the LLM is jointly optimized for response generation and unit test generation of different programming languages:
$\mathcal{L}_{all}=-\sum_{k=1}^{K}\mathbb{E}_{D^{L_{k}}}\log P(a^{L_{k}}|q^{L_{k}},c^{L_{k}})-\sum_{k=1}^{K}\mathbb{E}_{D^{L_{k}}}\log P(u^{L_{k}}_{v},u^{L_{k}}_{r}|q^{L_{k}},c^{L_{k}})$ (2)
where the responses from $D^{L_{k}}$ pass the unit tests of correctness and instruction-following. We can directly optimize the LLM by generating answers and unit tests.
3 Experiments
3.1 Experiment Setup
Evaluation Metric
To check whether the model-generated response is correct (‘Corr.’: Correctness Pass@111) and follows the user instructions (‘Instr.’: Instruction-following Pass@111), we evaluate LLMs by executing the generated code with unit tests (pass rate for just one-time generation) and separately report ‘Corr.’ and ‘Instr.’ Pass@111 score in tables.
Code LLMs
We report 40+ popular LLMs for IFEvalCode, including o1-mini (o1_mini, ), GPT4o (gpt4, ), Claude-3.5/3.7 (claude, ; claude37, ), Deepseek-v3/R1 (deepseekv3, ; deepseek_r1, ) llama3 (llama3, ), qwen3 (qwen, ), and code-specific models like Qwen2.5-Coder (qwencoder, ), DeepSeekCoder (deepseek_coder, ), CodeLlama (code_llama, ), and Codegemma (codegemma, ). Additionally, we fine-tune Qwen2.5-Coder-32B as our baseline ControlledCoder.
ControlledCoder Training Setup
Our constructed code generation dataset IFEvalCode-Instruct is for training ControlledCoder, ensuring fundamental instruction-following capabilities for code-related tasks. ControlledCoder, built on Qwen2.5-Coder-32B, is trained for 3 epochs using a cosine scheduler with an initial learning rate of $5\times 10^{-5}$ with a 3% warmup ratio. We employ AdamW (adamw, ) as the optimizer, with a batch size of 1024 and a maximum sequence length of 2048. (Details can be found in the Appendix).
Table 3: English evaluation of English queries in IFEvalCode of 8 programming languages. ‘Corr.’ denotes the correctness of the model-generated response, and ‘Instr.’ denotes the instruction-following accuracy of the response. The underlined values are the maximum in each column within each colored subgroup.
Models Params Python Java Cpp C-sharp Typescript Javascript Php Shell Avg.
Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr.
Closed-source LLMs
[…]
0.5B+ Open-source LLMs
Deepseek-Coder 1.3B 24.8 17.1 19.6 2.0 16.0 3.0 0.0 1.9 0.0 1.0 22.0 13.0 10.0 3.0 4.0 9.0 12.1 6.3
Qwen2.5-Coder 0.5B 18.1 19.0 8.8 2.0 8.0 2.0 9.7 1.9 2.0 18.0 13.0 12.0 11.0 4.0 3.0 9.0 9.3 8.5
Qwen2.5-Coder 1.5B 32.4 22.9 10.8 2.9 3.0 3.0 1.9 1.9 6.0 17.0 24.0 21.0 16.0 5.0 11.0 17.0 13.2 11.4
Qwen2.5-Coder 3B 35.2 24.8 11.8 2.0 11.0 5.0 11.7 1.9 0.0 18.0 26.0 29.0 22.0 12.0 15.0 23.0 16.7 14.4
Granite-Coder 3B 25.7 21.0 10.8 4.9 7.0 3.0 12.6 3.9 6.0 14.0 18.0 11.0 15.0 1.0 5.0 12.0 12.6 8.9
OpenCoder 1.5B 37.1 23.8 16.7 3.9 25.0 3.0 4.9 1.9 9.0 14.0 22.0 13.0 21.0 4.0 7.0 18.0 17.9 10.2
Yi-Coder 1.5B 32.4 21.0 13.7 2.0 13.0 4.0 11.7 2.9 8.0 17.0 26.0 21.0 16.0 4.0 8.0 14.0 16.2 10.7
Qwen3 0.6B 14.3 23.8 2.9 2.9 0.0 5.0 0.0 2.9 2.0 24.0 18.0 22.0 10.0 7.0 2.0 14.0 6.2 12.7
Qwen3-think 0.6B 13.3 15.2 2.0 2.0 1.0 5.0 0.0 1.9 2.0 6.0 8.0 3.0 13.0 0.0 0.0 6.0 4.9 4.9
Qwen3 1.7B 23.8 26.7 8.8 3.9 3.0 4.0 3.9 1.9 12.0 22.0 22.0 16.0 19.0 9.0 7.0 23.0 12.5 13.3
Qwen3-think 1.7B 25.7 21.0 3.9 2.9 3.0 2.0 0.0 2.9 10.0 11.0 22.0 17.0 28.0 2.0 7.0 13.0 12.5 9.0
Qwen3 4B 26.7 28.6 7.8 5.9 7.0 6.0 1.9 1.9 16.0 25.0 29.0 28.0 28.0 13.0 7.0 38.0 15.4 18
Qwen3-think 4B 34.3 21.9 12.7 4.9 6.0 4.0 1.0 3.9 12.0 12.0 29.0 20.0 25.0 2.0 13.0 17.0 16.7 10.7
6B+ Open-source LLMs
CodeLlama 7B 18.1 22.9 5.9 3.9 3.0 4.0 2.9 2.9 7.0 19.0 20.0 22.0 13.0 9.0 10.0 21.0 10.0 13.1
Llama3.1 8B 40.0 16.2 8.8 0.0 13.0 3.0 10.7 1.9 12.0 20.0 27.0 19.0 26.0 7.0 7.0 18.0 18.1 10.6
Deepseek-Coder 6.7B 37.1 21.9 27.5 7.8 20.0 3.0 4.9 1.9 16.0 17.0 30.0 20.0 30.0 8.0 12.0 19.0 22.2 12.3
Yi-Coder 9B 46.7 27.6 20.6 3.9 21.0 3.0 12.6 2.9 17.0 20.0 38.0 28.0 38.0 8.0 17.0 23.0 26.4 14.6
Granite-Coder 8B 35.2 19.0 11.8 3.9 11.0 3.0 3.9 1.9 13.0 15.0 32.0 15.0 21.0 1.0 11.0 8.0 17.4 8.4
OpenCoder 8B 38.1 24.8 21.6 5.9 24.0 4.0 5.8 1.9 16.0 17.0 42.0 19.0 26.0 9.0 14.0 15.0 23.5 12.1
CodeQwen1.5 7B 33.3 26.7 17.6 5.9 10.0 4.0 4.9 1.9 14.0 23.0 33.0 22.0 29.0 9.0 17.0 18.0 19.9 13.8
Qwen2.5-Coder 7B 41.9 27.6 25.5 3.9 16.0 3.0 11.7 2.9 19.0 22.0 36.0 30.0 43.0 11.0 20.0 28.0 26.7 16.0
Qwen3 8B 35.2 33.3 13.7 7.8 12.0 4.0 8.7 4.9 0.0 0.0 30.0 26.0 30.0 19.0 9.0 27.0 17.4 15.3
Qwen3-think 8B 35.2 31.4 15.7 9.8 6.0 4.0 1.9 5.8 16.0 14.0 30.0 24.0 28.0 2.0 22.0 14.0 19.4 13.2
14B+ Open-source LLMs
CodeLlama 13B 23.8 26.7 11.8 6.9 6.0 5.0 1.0 5.8 7.0 21.0 26.0 5.0 14.0 8.0 9.0 23.0 12.3 12.7
Qwen2.5-Coder 14B 48.6 27.6 27.5 7.8 22.0 2.0 18.4 2.9 20.0 25.0 38.0 31.0 34.0 12.0 20.0 35.0 28.6 17.9
Qwen3 14B 36.2 33.3 25.5 8.8 20.0 3.0 8.7 3.9 23.0 22.0 37.0 31.0 36.0 20.0 18.0 53.0 25.6 21.9
Qwen3-think 14B 51.4 31.4 22.5 15.7 16.0 5.0 1.9 4.9 30.0 24.0 48.0 33.0 43.0 4.0 24.0 17.0 29.6 16.9
Granite-Coder 20B 33.3 17.1 15.7 3.9 5.0 5.0 4.9 1.9 15.0 15.0 26.0 19.0 22.0 4.0 11.0 9.0 16.7 9.4
32B+ Open-source LLMs & MoE LLMs
Granite-Coder 34B 37.1 21.0 16.7 2.9 13.0 4.0 5.8 1.9 11.0 15.0 24.0 16.0 24.0 3.0 10.0 3.0 17.8 8.4
CodeLlama 34B 16.2 21.9 9.8 4.9 5.0 3.0 8.7 2.9 8.0 9.0 23.0 2.0 16.0 6.0 9.0 17.0 12.0 8.4
Deepseek-Coder 33B 41.9 22.9 33.3 4.9 22.0 4.0 7.8 1.9 24.0 20.0 33.0 18.0 37.0 7.0 19.0 20.0 27.3 12.3
Qwen2.5-Coder 32B 42.9 33.3 22.5 6.9 20.0 6.0 8.7 4.9 29.0 30.0 47.0 30.0 42.0 16.0 27.0 39.0 29.9 20.7
Llama3.1 70B 42.9 33.3 22.5 6.9 20.0 6.0 8.7 4.9 29.0 30.0 47.0 30.0 42.0 16.0 27.0 39.0 29.9 20.7
Qwen3 32B 49.5 33.3 25.5 8.8 17.0 7.0 10.7 1.9 20.0 29.0 38.0 34.0 41.0 28.0 22.0 52.0 28.0 24.2
Qwen3-think 32B 49.5 33.3 25.5 8.8 17.0 7.0 10.7 1.9 20.0 29.0 38.0 34.0 41.0 28.0 22.0 52.0 28.0 24.2
Qwen3 3/30B 54.3 30.5 30.4 11.8 31.0 4.0 1.0 4.9 31.0 26.0 48.0 28.0 46.0 2.0 36.0 21.0 34.7 16.0
Qwen3-think 3/30B 35.2 33.3 16.7 11.8 14.0 6.0 1.9 4.9 32.0 30.0 45.0 33.0 46.0 4.0 34.0 22.0 28.0 18.1
Qwen3 22B/235B 50.5 33.3 28.4 9.8 14.0 2.0 1.9 4.9 20.0 24.0 43.0 29.0 42.0 16.0 26.0 35.0 28.3 19.3
Qwen3-think 22B/235B 48.6 32.4 21.6 10.8 20.0 12.0 2.9 6.8 21.0 15.0 37.0 22.0 47.0 4.0 35.0 15.0 29.1 14.8
Deepseek-V3 37/637B 51.4 36.2 29.4 9.8 23.0 4.0 2.9 1.9 31.0 32.0 51.0 31.0 54.0 24.0 33.0 35.0 34.4 21.7
Deepseek-R1 37/637B 58.1 37.1 37.3 12.7 38.0 5.0 4.9 5.8 46.0 35.0 55.0 28.0 61.0 6.0 45.0 20.0 43.1 18.8
ControlledCoder 32B 58.3 38.2 39.3 13.5 30.0 12.0 10.7 6.8 44.0 38.0 47.0 38.0 48.0 30.0 45.0 52.0 40.3 34.2
Table 4: Chinese evaluation results in IFEvalCode of 8 programming languages. ‘Corr..’ denotes the correctness of the model-generated response, and ‘Instr.’ denotes the instruction-following accuracy of the response. The underlined fonts denote the best performance in the same parameter range.
Models Params Python Java Cpp C-sharp Typescript Javascript Php Shell Avg.
Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr. Corr. Instr.
Closed-source LLMs
[…]
0.5B+ Open-source LLMs
Deepseek-Coder 1.3B 22.9 10.5 14.7 1.0 7.0 2.0 1.9 1.9 0.0 1.0 16.0 13.0 7.0 7.0 1.0 2.0 8.9 4.8
Qwen2.5-Coder 0.5B 15.2 16.2 10.8 2.0 8.0 3.0 7.8 1.9 1.0 14.0 13.0 15.0 7.0 7.0 1.0 11.0 8.0 8.8
Qwen2.5-Coder 1.5B 28.6 21.9 9.8 2.9 2.0 5.0 0.0 1.9 6.0 20.0 28.0 23.0 18.0 11.0 4.0 24.0 12.1 13.7
Qwen2.5-Coder 3B 31.4 20.0 13.7 3.9 12.0 5.0 11.7 2.9 7.0 19.0 31.0 23.0 27.0 12.0 10.0 30.0 18.0 14.4
Granite-Coder 3B 25.7 21.0 10.8 4.9 7.0 3.0 12.6 3.9 6.0 14.0 18.0 11.0 15.0 1.0 5.0 12.0 12.6 8.9
OpenCoder 1.5B 39.0 16.2 21.6 3.9 20.0 2.0 6.8 1.9 10.0 14.0 18.0 14.0 21.0 8.0 7.0 17.0 18.0 9.6
Yi-Coder 1.5B 34.3 17.1 15.7 1.0 16.0 5.0 14.6 1.9 10.0 11.0 22.0 21.0 15.0 7.0 2.0 13.0 16.3 9.6
Qwen3 0.6B 11.4 21.0 2.9 3.9 1.0 6.0 0.0 1.9 0.0 25.0 13.0 19.0 8.0 15.0 0.0 22.0 4.6 14.2
Qwen3-think 0.6B 10.5 10.5 2.9 3.9 0.0 2.0 0.0 1.9 1.0 10.0 9.0 8.0 6.0 1.0 0.0 7.0 3.7 5.6
Qwen3 1.7B 27.6 29.5 6.9 3.9 4.0 5.0 1.0 1.9 9.0 25.0 19.0 19.0 15.0 13.0 5.0 31.0 11.0 16.0
Qwen3-think 1.7B 21.0 14.3 2.9 2.9 0.0 0.0 0.0 1.0 10.0 1.0 14.0 3.0 21.0 2.0 3.0 3.0 9.0 3.5
Qwen3 4B 24.8 30.5 6.9 7.8 5.0 4.0 2.9 3.9 14.0 26.0 28.0 31.0 29.0 20.0 8.0 47.0 14.8 21.2
Qwen3-think 4B 24.8 7.6 4.9 1.0 1.0 1.0 0.0 0.0 9.0 5.0 16.0 3.0 21.0 1.0 14.0 6.0 11.4 3.1
6B+ Open-source LLMs
CodeLlama 7B 0.0 1.9 5.9 1.0 1.0 3.0 0.0 2.9 5.0 2.0 23.0 0.0 1.0 2.0 2.0 17.0 4.7 3.7
Llama3.1 8B 35.2 21.9 9.8 2.0 13.0 2.0 15.5 1.9 10.0 19.0 24.0 23.0 24.0 14.0 5.0 24.0 17.2 13.5
Deepseek-Coder 6.7B 38.1 23.8 20.6 3.9 16.0 3.0 2.9 1.9 18.0 16.0 34.0 17.0 25.0 10.0 16.0 23.0 21.4 12.3
Yi-Coder 9B 43.8 24.8 31.4 5.9 21.0 2.0 7.8 1.9 18.0 16.0 33.0 24.0 46.0 11.0 14.0 23.0 26.9 13.6
Granite-Coder 8B 32.4 19.0 11.8 3.9 14.0 4.0 6.8 1.9 9.0 13.0 29.0 19.0 9.0 3.0 10.0 16.0 15.3 10.0
OpenCoder 8B 40.0 25.7 17.6 4.9 16.0 5.0 4.9 1.9 15.0 19.0 30.0 19.0 32.0 12.0 13.0 26.0 21.1 14.2
CodeQwen1.5 7B 41.0 21.9 15.7 4.9 12.0 4.0 9.7 1.9 12.0 23.0 33.0 25.0 27.0 11.0 13.0 23.0 20.5 14.3
Qwen2.5-Coder 7B 42.9 27.6 21.6 4.9 11.0 5.0 6.8 2.9 16.0 21.0 39.0 29.0 36.0 11.0 21.0 32.0 24.3 16.7
Qwen3 8B 36.2 32.4 11.8 5.9 7.0 5.0 5.8 3.9 0.0 0.0 29.0 32.0 32.0 27.0 6.0 18.0 16.0 15.6
Qwen3-think 8B 28.6 12.4 8.8 2.9 3.0 1.0 1.9 1.0 14.0 2.0 26.0 6.0 27.0 0.0 15.0 6.0 15.6 4.0
14B+ Open-source LLMs
CodeLlama 13B 19.0 22.9 10.8 3.9 7.0 4.0 4.9 2.9 7.0 24.0 24.0 8.0 22.0 7.0 4.0 22.0 12.3 11.9
Qwen2.5-Coder 14B 47.6 29.5 26.5 8.8 22.0 3.0 19.4 2.9 24.0 32.0 41.0 30.0 39.0 15.0 19.0 36.0 29.9 19.6
Qwen3 14B 39.0 32.4 19.6 8.8 17.0 3.0 10.7 4.9 23.0 31.0 41.0 33.0 40.0 23.0 18.0 53.0 26.0 23.6
Qwen3-think 14B 49.5 26.7 16.7 2.9 10.0 1.0 2.9 1.0 23.0 8.0 34.0 6.0 37.0 1.0 22.0 10.0 24.4 7.2
Granite-Coder 20B 30.5 21.9 15.7 2.9 4.0 4.0 6.8 3.9 13.0 13.0 25.0 19.0 26.0 3.0 4.0 10.0 15.7 9.8
32B+ Open-source LLMs & MoE LLMs
Granite-Coder 34B 34.3 23.8 12.7 4.9 10.0 4.0 5.8 1.9 16.0 15.0 29.0 23.0 24.0 4.0 5.0 9.0 17.2 10.7
CodeLlama 34B 4.8 15.2 5.9 2.0 15.0 3.0 12.6 1.9 10.0 6.0 18.0 0.0 17.0 10.0 3.0 22.0 10.7 7.5
Deepseek-Coder 33B 44.8 21.0 26.5 4.9 24.0 7.0 7.8 1.9 21.0 18.0 35.0 20.0 30.0 14.0 20.0 26.0 26.2 14.1
Qwen2.5-Coder 32B 43.8 34.3 28.4 8.8 19.0 2.0 12.6 5.8 25.0 29.0 41.0 34.0 44.0 24.0 23.0 38.0 29.6 22.0
Llama3.1 70B 47.6 22.9 14.7 2.0 19.0 3.0 11.7 2.9 19.0 24.0 35.0 28.0 39.0 19.0 21.0 40.0 25.9 17.7
Qwen3 32B 42.9 41.0 24.5 7.8 17.0 7.0 9.7 3.9 27.0 25.0 38.0 31.0 40.0 23.0 19.0 51.0 27.3 23.7
Qwen3-think 32B 54.3 19.0 30.4 5.9 24.0 2.0 7.8 1.9 26.0 2.0 42.0 7.0 50.0 1.0 30.0 6.0 33.1 5.7
Qwen3 3/30B 43.8 33.3 28.4 11.8 19.0 6.0 5.8 5.8 24.0 31.0 42.0 31.0 38.0 29.0 21.0 48.0 27.8 24.4
Qwen3-think 3/30B 42.9 21.0 16.7 2.9 12.0 0.0 2.9 2.9 21.0 4.0 41.0 10.0 40.0 1.0 22.0 7.0 24.8 6.2
Qwen3 22B/235B 46.7 38.1 22.5 8.8 12.0 3.0 1.0 2.9 25.0 24.0 44.0 30.0 47.0 22.0 20.0 38.0 27.3 20.9
Qwen3-think 22B/235B 46.7 11.4 26.5 2.0 26.0 1.0 1.0 0.0 32.0 1.0 43.0 2.0 43.0 1.0 36.0 2.0 31.7 2.6
Deepseek-V3 37/671B 50.5 35.2 32.4 9.8 21.0 5.0 4.9 1.9 32.0 34.0 51.0 25.0 51.0 29.0 34.0 38.0 34.6 22.2
Deepseek-R1 37/671B 58.1 37.1 37.3 12.7 38.0 5.0 4.9 5.8 46.0 35.0 55.0 28.0 61.0 6.0 45.0 20.0 43.1 18.8
ControlledCoder 32B 54.8 39.2 40.3 13.8 34.0 7.0 11.7 6.8 48.0 35.0 49.0 40.0 48.0 32.0 51.0 55.0 42.1 28.6
3.2 Main Results
Results of English queries in IFEvalCode
Table 3 presents evaluation results of “Correctness” (Corr.) and “Instruction Compliance” (Instr.) for 40+ LLMs across 888 programming languages (Python, Java, C++, C#, TypeScript, JavaScript, PHP, Shell). These models vary in size and origin (closedsource/opensource/mixed expert models) and are tested on English programming queries. Closed-source LLMs (such as Claude-3.7, GPT-4.1, and various mini versions) generally achieve overall correctness rates concentrated between 30% and 40%. Instruction compliance rates typically hover around 20% to 25%. Open-source LLMs with 0.5B to 3B parameters generally demonstrate limited performance, with average correctness rates mostly between 6% and 18%. The top performer in this range, OpenCoder (1.5B), only reaches 17.9%. This indicates that small-scale models still struggle to achieve satisfactory results in cross-language, multi-task coding scenarios. Open-source models in the 6B to 14B parameter range show significant improvement. The 7B Qwen2.5-Coder, 8B Qwen3-think, and 14B Qwen3 models all surpass average correctness rates of 25% to 30%. Notably, the 14B Qwen3-think achieves a multilingual overall correctness rate of 29.6%. Among models with 32B parameters or more / MoE architectures, Deepseek-R1 (37/637B) performs the best, reaching an average correctness rate as high as 43.1%. It sets the highest group scores across multiple languages: Python (58.1%), Java (37.3%), C++ (38.0%), TypeScript (46.0%), JavaScript (55.0%), PHP (61.0%), and Shell (45.0%). Its instruction compliance rate also improves to around 19% to 20%. Overall, model scale and architectural complexity (such as MoE expert models) significantly boost programming task capabilities. At the same time, whether closed-source or open-source, the current instruction compliance rates remain noticeably lower than correctness rates. This highlights the need for further optimization in instruction tuning and code execution verification in the future.
Results of Chinese queries in IFEvalCode
Table 4 presents the evaluation results of various large models on Chinese programming queries (covering eight languages: Python, Java, C++, C#, TypeScript, JavaScript, PHP, and Shell) in terms of "Correctness" (Corr.) and "Instruction Compliance" (Instr.). The best performers within each size range are highlighted in bold. Closed-source models (Claude-3.x, GPT-4.0/4.1, and their mini versions) have an overall average correctness between 27% and 38%, with instruction compliance around 20% to 25%. Among them, GPT-4.1 leads the closed-source group in both correctness (36.4%) and instruction compliance (24.7%). Although o1-mini and o3-mini show higher correctness rates (39.3% and 37.8%, respectively), their instruction compliance rates are relatively lower (23.5% and 24.1%). Open-source models in the 0.5B–3B parameter range show limited performance, with average correctness mostly between 5% and 18%. OpenCoder (1.5B) at 18.0% and Qwen2.5-Coder (3B) also at 18.0% lead this group, indicating that small-scale models still need improvement in multilingual coding tasks. In the 6B–14B parameter range, open-source models begin to show significant improvement—Yi-Coder (9B) achieves an average correctness of 26.9%, while Qwen2.5-Coder (14B) ranks first in this group with a correctness of 29.9% and instruction compliance of 19.6%. Qwen3 (8B) also surpasses a 23% correctness rate and obtains 32% instruction compliance on Python and TypeScript. Among large models with 32B+ parameters and MoE (Mixture of Experts) architectures, Deepseek-R1 (37/671B) demonstrates the most outstanding performance, with an average correctness of 43.1% and the highest scores across multiple sub-tasks including Python, Java, C++, TypeScript, JavaScript, PHP, and Shell. Although its instruction compliance (18.8%) is not the highest, it remains relatively high. Deepseek-V3 follows closely (34.6%), with Qwen2.5-Coder 32B (29.6%) and Llama3.1-70B (25.9%) ranking next. Overall, model scale and architecture (especially MoE expert models) have a significant positive impact on Chinese programming capabilities. Meanwhile, instruction compliance rates are generally lower than correctness rates across all models, highlighting the need for further optimization in instruction tuning and multilingual semantic understanding in the future.
4 Further Analysis
Ablation Study.
Table 5: Ablation study.
ID Method Corr.avg Instr.avg
Qwen2.5-Coder-32B 29.9 20.7
+ $D_{s_{1}}$ 34.5 26.4
+ $D_{s_{2}}$ 38.2 30.4
+ Iter2 40.3 34.2
+ Iter3 40.5 33.2
In Table 5, we conduct the ablation study to explore the effect of each component in IFEvalCode-Instruct. ControlledCoder is fine-tuned on the base Qwen2.5-Coder by combining the instruction corpus from the forward and backward constraints generation ( and outperform the baseline Qwen2.5-Coder ). The fine-tuned coder ControlledCoder can iteratively update the instruction corpus IFEvalCode-Instruct using forward and backward constraints generation. We can see that joint multilingual rejecting fine-tuning can effectively bring improvement for controlled code generation.
(a) Distribution of constraints number
(b) Scores of different constraints number
Figure 5: Figure 5(a) and 5(b) plot the trends with the number of constraints increaing
Correlation between Correctness and Code Instruction Following
Figure 6 displays the correctness overlap between correctness score (Eh), instruction following score (Zh), correctness score (En), and instruction following score (Zn) of 888 different programming languages.
Figure 6: The Venn diagram of the numbers of the correct samples for each programming language.
Figure 6 shows a grid of four-set Venn diagrams—one subplot per programming language—where each diagram partitions the model’s outputs into the following four evaluation sets: (1) Chinese-prompt correctness, (2) Chinese-prompt instruction compliance, (3) English-prompt correctness, and (4) English-prompt instruction compliance. By inspecting the overlaps, we can immediately see several consistent patterns across all eight languages: Correctness dominates. Both the Chinese-correctness and English-correctness sets are substantially larger than their corresponding instruction-compliance sets, indicating that LLMs more often produce functionally correct code than code that also respects the extra constraints. Cross-language consistency in correctness. The overlap between Chinese-correctness and English-correctness is very large in every language, showing that whether the prompt is given in Chinese or English has little effect on a model’s ability to generate a passing solution. Instruction compliance is more fragile. The Chinese-instruction and English-instruction sets are much smaller, and their overlap is modest. In other words, even if a model is correct in both languages, it frequently fails to satisfy the same set of style or structural constraints when the prompt language switches. The smallest region is the four-way intersection. Only a small fraction of examples are simultaneously correct and compliant in both Chinese and English, demonstrating that achieving full controlled generation across languages remains the hardest challenge.
5 Related Work
Code Large Language Model
The rapid progress of large language models(gpt4, ; llama2, ; llama3, ; deepseekv3, ; deepseek_r1, ) has enabled complex code-related tasks. Early models (e.g. BERT (bert, ) and GPT (gpt, )) trained on billions of code snippets, focused on code understanding and generation (codex, ; code_bert, ; bloom, ; AlphaCode, ; codet5, ; santacoder, ). Recent advances in domain-specific pre-training and instruction fine-tuning (kun, ; mammoth2, ) have enhanced models like Qwen2.5-Coder (qwen_coder, ) and Deepseek-Coder (deepseek_coder, ), achieving strong performance in code completion, synthesis, and repair.
Code Benchmarks
In the field of code, a rich tapestry of benchmarks (codex, ; mbpp, ; humaneval_x, ; codereval, ; arcade_nl2code, ; humaneval_xl, ; xcodeeval, ; BabelCode, ) has been developed to address the challenges of assessing code quality, functionality, and efficiency. Current benchmarks support the evaluation of various tasks, including code generation (e.g. NCB (naturalcodebench, ) and MultiPL-E (multipl_e, )), code repair  (e.g. DebugBench (debugbench, ) and MdEval (mdeval, )), code competition (e.g. LCB (livecodebench, ) and CodeElo (codeelo, )), software development (e.g. Aider and SWE-Bench (swe_bench, )). But existing code benchmarks ignore the multilingual evaluation of controlled code generation for LLMs. IFEval (ifeval, ) emphasizes the importance of evaluating the proficiency of LLMs in instruction following, where IFEval contains a series of “verifiable instructions”.
6 Conclusion
In this paper, we first introduce IFEvalCode, a multilingual benchmark with 1.6K samples across 8 programming languages, designed to evaluate both code correctness and code controllability in LLM-generated code. To address these challenges, we propose a novel framework that combines forward constraints generation and backward generation, enhancing LLMs’ ability to follow human instructions for controlled code generation. Our approach advances the reliability and precision of LLMs in adhering to user intent while producing functionally accurate code.
Appendix A List of Large Language Models
Table 6 lists all the large language models evaluated in this study and their download links; we then systematically evaluate the performance of these models on the IFEvalCode benchmark.
Table 6: Evaluation Model List
Model Download Link
Claude-3.7-Sonnet https://www.anthropic.com/news/claude-3-7-sonnet
Claude-3.5-Sonnet https://www.anthropic.com/news/claude-3-5-sonnet
GPT-4o https://platform.openai.com/docs/models/gpt-4o
GPT-4o-mini https://platform.openai.com/docs/models/gpt-4o-mini
GPT-4.1 https://platform.openai.com/docs/models/gpt-4.1
GPT-4.1-mini https://platform.openai.com/docs/models/gpt-4.1-mini
o1-mini https://platform.openai.com/docs/models/o1-mini
o3-mini https://platform.openai.com/docs/models/o3-mini
o4-mini https://platform.openai.com/docs/models/o4-mini
grok-3 https://docs.x.ai/docs/models#models-and-pricing
grok-3-mini-fast https://docs.x.ai/docs/models#models-and-pricing
Deepseek-Coder-1.3B https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-instruct
Qwen2.5-Coder-0.5B https://huggingface.co/Qwen/Qwen2.5-Coder-0.5B
Qwen2.5-Coder-1.5B https://huggingface.co/Qwen/Qwen2.5-Coder-1.5B
Qwen2.5-Coder-3B https://huggingface.co/Qwen/Qwen2.5-Coder-3B
Granite-Coder-3B https://huggingface.co/ibm-granite/granite-3b-code-base-128k
OpenCoder-1.5B https://huggingface.co/infly/OpenCoder-1.5B-Instruct
Qwen3-0.6B https://huggingface.co/Qwen/Qwen3-0.6B
Qwen3-1.7B https://huggingface.co/Qwen/Qwen3-1.7B
Qwen3-4B https://huggingface.co/Qwen/Qwen3-4B
CodeLlama-7B https://huggingface.co/meta-llama/CodeLlama-7b-Instruct-hf
Deepseek-Coder-6.7B https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct
Yi-Coder-9B https://huggingface.co/01-ai/Yi-Coder-9B-Chat
Granite-Coder-8B https://huggingface.co/ibm-granite/granite-8b-code-instruct-128k
OpenCoder-8B https://huggingface.co/infly/OpenCoder-8B-Instruct
CoderQwen1.5-7B https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat
Qwen2.5-Coder-7B https://huggingface.co/Qwen/Qwen2.5-Coder-7B
Qwen3-8B https://huggingface.co/Qwen/Qwen3-8B
Qwen2.5-Coder-14B https://huggingface.co/Qwen/Qwen2.5-Coder-14B
Qwen3-14B https://huggingface.co/Qwen/Qwen3-14B
Granite-Coder-20B https://huggingface.co/ibm-granite/granite-20b-code-instruct-8k
Granite-Coder-34B https://huggingface.co/ibm-granite/granite-34b-code-instruct-8k
Deepseek-Coder-33B https://huggingface.co/deepseek-ai/deepseek-coder-33b-instruct
Qwen2.5-Coder-32B https://huggingface.co/Qwen/Qwen2.5-Coder-32B
Qwen3-32B https://huggingface.co/Qwen/Qwen3-32B
Qwen3-30B-A3B https://huggingface.co/Qwen/Qwen3-30B-A3B
Qwen3-235B-A22B https://huggingface.co/Qwen/Qwen3-235B-A22B
DeepSeek-V3 https://huggingface.co/deepseek-ai/DeepSeek-V3
DeepSeek-R1 https://huggingface.co/deepseek-ai/DeepSeek-R1
Appendix B Limitations
Programming Languages
To ensure a more comprehensive evaluation of multilingual controlled code generation, it would be beneficial to expand IFEvalCode to include additional programming languages, particularly those that are less commonly used or serve niche applications, as the current coverage of eight languages may not fully represent the diversity and specialized use cases present in real-world software development.
Complex Controlled Code Generation
While IFEvalCode aims to simulate realistic controlled code generation scenarios, IFEvalCode ignores the complex controlled code generation and lacks repository-based evaluation.
Appendix C Ethical Considerations
C.1 Potential Risks
IFEvalCode serves as a comprehensive assessment tool for controlled code generation. However, improper use may lead to risks such as incorrect program analysis, faulty repairs, or even system failures. To mitigate these risks, we strongly recommend conducting evaluations within a sandbox environment, which effectively isolates potential hazards and safeguards both the accuracy of the assessment and the stability of the underlying system.
Appendix D Human Annotation
Annotators
To construct IFEvalCode, we developed a rigorous and systematic human annotation process, guided by meticulously designed annotation protocols, to ensure the accuracy, consistency, and high quality of questions. This process integrates multiple quality control measures at every stage to maintain reliability and precision across all collected data. We recruited 666 programmers with computer science degrees equipped with strong computer science fundamentals.
Appendix E Analysis of Instruction Constraints
Illustrated in Figure 2(b), data annotation was based on diverse, real-world constraint types frequently encountered in software engineering. In our annotation, key challenge clusters for LLMs are:
  • Naming constraints (e.g., “use camelCase” or “no abbreviation”): Models often ignore subtle naming conventions unless strongly prompted.
  • Syntax restrictions (e.g., “no recursion”, “use list comprehension”): Many LLMs hallucinate or use default constructs, especially in multi-turn or long/intricate tasks.
  • Variable constraints (e.g., variable reuse/uniqueness): LLMs may create unnecessary variables or reuse names incorrectly.
[…]
The most challenging for LLMs are “style constraints,” “naming constraints,” and “performance/resource constraints,” as these are not always explicit in training corpora and often conflict with models’ default generations. Performance varies by LLM family; for example, Claude Series models are slightly better, but none are consistently reliable across all constraint categories.
Appendix F Related Work
Code Large Language Models.
Large language models (LLMs) designed for coding tasks have demonstrated exceptional capabilities in code generation, debugging, translation, and other essential functions for modern software engineering [14, 4, 43, 57], with numerous benchmarks developed to evaluate these capabilities, though many focus on a limited selection of programming languages like Python and Java [82, 6, 28]. Recent advancements in code LLMs, such as Code Llama [53], DeepSeek-Coder [22], OpenCoder [23], and Qwen2.5-Coder [26], have significantly improved multilingual code generation and debugging, evaluated using benchmarks like MultiPL-E [10], McEval [11], and MdEval [39]. Code generation, a fundamental task for code LLMs, requires interpreting natural language descriptions and generating corresponding code snippets [20, 31, 37, 34], with benchmarks proposed for code translation [68], code completion [8, 77], and structured data understanding [65, 56].
Instruction Following.
Automatic program debugging holds substantial practical value. With the emergence of LLM capabilities, a growing number of individuals are utilizing LLMs for code debugging, leading to extensive research in this field. Code Debugging includes several tasks such as bug or vulnerability detection [47, 3, 74, 76, 84], fuzz test [15, 66, 70], program repair [64, 35, 78, 49, 19, 58, 63], GitHub issues auto resolving [29, 12, 59]. To effectively assess the code debugging capabilities of LLMs, several benchmark tests have been introduced [48, 55, 67, 78, 60, 71]. Notably, DebugBench [60] provides a comprehensive classification of error types and analyzes the debugging capabilities of LLMs based on these categories. Similarly, DebugEval [71] has designed various debugging-related tasks to evaluate LLM performance across different task dimensions. However, these studies focus on 1 to 3 languages. In reality, there are significant differences in code errors between languages, leading to numerous language-specific errors. To address this gap, we propose IFEvalCode, a comprehensive code debugging benchmark covering 20 languages, aiming to assess LLM debugging capabilities from a broader perspective.